# Optimal Spatiotemporal Resolutions

### v2

@@ Results
It is well-known that complex systems can be governed by local or collective events [cioni_innate_2023, crippa_detecting_2023, caruso_timesoap_2023, de_marco_controlling_2021, butail_model-free_2016, offei-danso_high-dimensional_2022] that may emerge and amplify within them. When one does not have prior knowledge of a system, choosing the best analysis setup to capture the effects/phenomena characterizing its behavior may be non-trivial. Nonetheless, such a choice can influence or bias the analysis, potentially revealing or obscuring distinct types of information, depending on the underlying nature and physics of the system (Fig. 1F), as we are going to discuss in the subsequent analyses.
In the case of aqueous systems, it is known that focusing solely on the first solvation shell can be limiting [crippa_detecting_2023, offei-danso_high-dimensional_2022]. It has been recently demonstrated how analyses extending beyond that (e.g., including up to the fifth nearest neighbor)[martino2024data], can offer a more comprehensive understanding of the system [donkor_machine-learning_2023, donkor_beyond_2024, foffi_correlated_2023, soper_is_2019, daidone_statistical_2023]. This raises important questions such as, e.g.: What is the effect of the spatial resolution on the results that can be effectively attained upon analysis? Is there an “optimal resolution” to analyze a system? And is it possible to somehow determine it a priori? Addressing such questions is challenging, but at the same time particularly useful, especially in the study of systems about which little is known a priori.
Fig. 1 shows the results of analyses obtained calculating the LENS time-series using a cutoff of $r_{c}=10$. While an $r_{c}$ around such value is commonly assumed to enable information-rich analyses of aqueous systems, [crippa_detecting_2023, martino2024data, caruso_timesoap_2023, donkor_machine-learning_2023, Capelli2022Ephemeral] we introduce an automated approach to systematically determine the optimal spatial resolution to study virtually any type of system. It is important to emphasize that the extractable information – and the corresponding optimal cutoff – depends both the internal physics of the system and the descriptor’s effectiveness in capturing relevant physical features. The efficiency and signal-to-noise ratio are in fact intrinsic properties of each descriptor [martino2024data]. Here, using LENS as an example, we present a systematic, data-driven approach to determine the optimal resolution for studying the internal dynamic complexity and overall physics of the system.
@@
We calculated the LENS descriptor for all oxygen atoms throughout the trajectory, using the identified key distances as cutoffs. For each cutoff, our method applied Onion Clustering to analyze the time-series and determine the number of statistically distinct environments detected. By gathering and comparing these results, we identified the “optimal spatiotemporal resolution” for extracting physically relevant information from the trajectories. At this stage, the analysis depends solely on the chosen descriptor.
@@
Fig. 2B-G show the results obtained via Onion Clustering of the LENS time-series computed with the different cutoffs. Fig. 2B illustrates the results of Onion Clustering applied to LENS time-series computed with $r_{c}=3.3$, corresponding to the first solvation shell. At all time resolutions, the analysis detects at most one dynamically persistent environment, the bulk (ice), while all other molecular data points are grouped into the unclassifiable data cluster (transitions). Even at the highest resolution, the fraction of unclassifiable data points is approximately 50%, highlighting the limited effectiveness of considering only the first solvation shell for this aqueous system. This outcome is akin to conventional pattern recognition methods. Moreover, for time resolutions below $\Delta t>10$ ns, Onion Clustering fails to resolve any dynamically persistent clusters (Fig. 2B), as the fraction of unclassifiable data points (orange) reaches 100%. These results show that restricting the analysis to the first solvation shell does not includes in the LENS descriptor enough information for a meaningful classification of the system’s dynamics.
@@
Notably, when the second solvation shell is included using a cutoff of $r_{c}=7.5$ (Fig. 2C), Onion Clustering identifies 2 to 3 dynamically persistent environments. Crucially, the ice/water interface becomes discernible within this time-series for resolutions between 2 and 8 ns. This observation aligns with findings reported for other descriptors [martino2024data, lionello2024relevant], which highlight the existence of a lower time resolution threshold necessary to identify certain environments. In fact, for $\Delta t>8$ ns, the analysis loses the ability to distinguish the ice/water interface as a separate environment. Interestingly, Fig. 2C also shows that at very high time resolutions ($\Delta t<2$ ns), information about the ice/water interface is again lost, reducing the number of detectable clusters to two. This, as recently discussed  [martino2024data, lionello2024relevant], is due to the fact that there is a minimum observation time which is required so that transitions can be robustly distinguished from the more dominant states, and thus classified as separate environments.
@@
When the cutoff distance is increased beyond 8.5 Å (Fig. 2D-G), including the third solvation shell, the LENS time-series consistently resolve three dynamically persistent clusters down to a resolution of $\Delta t\sim 10$ ns. This clustering stability persists up to a cutoff distance of approximately 15 Å. In other words, spatial resolutions in the range of $8.5$to 15 Åoffer the best performance, as they maximize the detection of distinct clusters (Fig. 2D-F) and minimize information loss (orange curve) for time resolutions $\Delta t<10$ ns. For this system, these spatial resolutions, when applied to the analysis of LENS time-series, enable the extraction and representation of the system’s dominant features with optimal efficiency.
@@
Figure 3: Effect of spatial and temporal resolutions on information extraction from, e.g., SOAP time-series data. A Schematic representation of SOAP. For each molecule $i$ in the system, at each sampled timestep $t$, the SOAP vector contains information on the distances and spatial displacements of the neighbors within a sphere of cutoff $r_{c}$: the SOAP power spectrum is thus a fingerprint of the local neighbors density around every SOAP center (the oxygen atoms of each water molecule, in this case). B-G Onion plots reporting the results of the clustering analyses of the SOAP PC1 time-series as a function of the cutoff $r_{c}$ values used to calculate the SOAP power spectra. H In green: mean number of classifiable micro-clusters (ENVs) before the fraction of unclassifiable data reaches 50%, as a function of the $r_{c}$ used in the analysis. The data show a clear trend (green curve), where the maximum efficiency in information extraction-and-discretization is obtained at $13\leq r_{c}\leq 15$, where the interface is detected in a robust way down to time-resolutions of $\Delta t<20$ ns.
@@
When the LENS cutoff radius is set to values beyond $15$, the effectiveness of the analysis begins to diminish. We extended the tests up to the maximum cutoff of 30Å, constrained by the size of the simulation box. As shown in Fig. 2B-G and in supplementary Figures S1, the ability to distinguish the ice/water interface progressively deteriorates with increasing cutoffs. At higher values, the analysis yields results similar to those obtained with smaller cutoffs, where only the solid ice phase can be classified as a stable environment.
@@
This result provides critical insights into the physics of the system, revealing the characteristic scales of the environments dominating its dynamics. In this case, this ice/water system exhibits internal dynamics driven by collective events with specific spatial and temporal scales. For instance, the ice/water interface has a characteristic thickness of around 1.5 nm and contains molecules with a residence time of up to $\sim 10$ ns (Fig. 2I). Clearly, this “optimal spatiotemporal resolution” depends on the dynamics of each system (see next paragraph) and on the capability of different descriptors to capture them. Each descriptor emphasizes distinct features and has its own signal-to-noise ratio [martino2024data]. To illustrate the generality of these concepts, we repeated the Onion Clustering analyses on the first principal component (PC1) time-series of the SOAP spectra computed for each molecule, calculated with different cutoff radii as done for LENS. This serves as another example of a descriptor providing rich information on the system (see Method for details).
@@
The results shown in Fig. 3 show a very similar trend as those in Fig. 2. As in the LENS analysis, Fig. 3A-G demonstrates the existence of an optimal spatial resolution that maximizes the information extracted from the SOAP PC1 time-series. Below this resolution, the analysis lacks sensitivity, and above it, the time-series become too noisy to resolve the ice/water interface. All other explored spatial resolutions are shown in supplementary Figure S2. Also MD representative snapshot of this system, all taken at $\Delta t=40$ ns are shown in supplementary Figure S5. The plot in Fig. 3H shows a trend similar to Fig. 2H, revealing an optimal spatial resolution at $r_{c}=13-15$. With this $r_{c}$, the interface is detectable using a time-resolution $\Delta t<20$ ns (Fig. 3D-E). Interestingly, the slightly higher optimal cutoff radius for SOAP compared to LENS reflects differences between the descriptors: while LENS captures exchanges between molecules inside and outside the cutoff, SOAP considers only the local environment within the cutoff. This makes the SOAP optimal radius consistent with the detection of events at the fourth solvation shell, compared to the third shell for LENS, while capturing the same underlying physical phenomena. Let’s now apply these concepts to another system with completely different dynamical complexity to investigate how the approach performs in distinct scenarios.
@@
Figure 4: Optimal resolution in the study of a dynamic metal surface. A MD snapshots of a Cu surface composed of 2400 atoms simulated at $T=600$ K: the atoms are colored according to their coordination number. Top: crystalline surface at $T=0$ K before simulation start; bottom: representative snapshot of the surface equilibrated at $T=600$ K. B Onion plot: number of clusters (in blue) classified from the LENS time-series obtained with cutoff radius $r_{c}=5$, and fraction of unclassifiable data points (orange) as a function of the time-resolution $\Delta t$ used in the analysis. The vertical red dashed line indicates $\Delta t=0.12$ ns as an example time-resolution used to plot the analysis results in panels C-D. C LENS time-series for all Cu atoms in the surface model along the entire MD simulation of $\tau=150$ ns (sampled every $\Delta\tau=10$ ps - see Methods for complete details). At the resolution of $\Delta t=0.12$ ns, Onion Clustering classifies six different micro-clusters (colored areas) whose thresholds are identified by horizontal colored dashed lines. D Top view of the simulation box at three different simulation times: $t=40$, $61$, and $150$ ns. An example Cu atom (ID = 144) is highlighted in black, whose history trajectory up to that point is colored according to the LENS micro-cluster it belonged to. At 150 ns, a red/orange straight vertical sliding on the (211) surface edge is clearly visible. E Radial distribution function $g(r)$ of the Cu atoms in the system: relevant spatial cutoffs are highlighted by red circles (relevant $g(r)$ minima). F-M Onion plots reporting the results of the analyses on LENS time-series with different $r_{c}$. Same color coding as before. N Mean number of clusters (ENVs) classifiable (before ENV0 reaches 50%) as a function of the $r_{c}$ used in the analysis. In this case, dominated by local single-atom dynamical events, the optimal resolution is achieved at the smallest $r_{c}$: i.e., when accounting only for the first neighbors shell.
@@
To demonstrate the generality of our approach, we now apply it to analyze LENS time-series computed from simulation of an atomistic model of a copper (211) surface simulated at 600 K (Fig. 4A) composed by 2400 Cu atoms replicating on the $xy$ plane via periodic boundary conditions, using a deep neural network potential trained to guarantee DFT-level accuracy (see Method section for details) [cioni_innate_2023]. As recently demonstrated, even well below the melting temperature $\sim 1400$ K, such system demonstrates a peculiar dynamics, with sparse individual atoms migrating along the surface crests and diffusing sliding rapidly over the surfaces face, being then reincorporated into the crests [cioni_innate_2023, crippa_detecting_2023, crippa_machine_2023, caruso2024classification]. These rare local events are challenging to capture using conventional pattern recognition method, while they can be effectively captured via Onion Clustering analyses of LENS time-series [becchi_layer-by-layer_2024].
Using this MD trajectory, we followed the same procedure illustrated in the previous paragraph, computing LENS using different cutoff radii (using different $g(r)$ minima, Fig. 4E) and then measuring, via Onion Clustering, how much information can be extracted for each time-series dataset. As shown in Fig. 4, Onion Clustering can detect different environments. For every choice of the cutoff radius, the clustering can identify various, more or less dynamic, environments in the system (Fig. 4B-C, for instance, using $r_{c}=5$) corresponding, e.g., to the metal bulk or static surface domains, as well as the crests and diffusing atoms on the surface (Fig. 4D). The trend in the number of resolvable clusters (Fig. 4F-M: in blue) shows how the number of distinct environments that Onion Clustering can resolve in the system decreases as the cutoff radius used in the analysis increases (Fig. 3B,F-M). In Fig. 4F, for example, it is evident that up to 8-10 clusters can be resolved using $r_{c}=3$(first neighbors shell) up to a temporal resolution of $\Delta t\sim 100$ ps (the faster dynamics occurring in this system require an higher temporal resolution than the ice/water system studied previously). As the cutoff increases, the number of clusters tends to decrease, reaching a maximum of 8 for a cutoff radius in the range $6.5\leq r_{c}\leq 10.5$. For $r_{c}>14$, there is a considerable drop in the number of resolved clusters, and above $r_{c}=15$the effectiveness of the analysis deteriorates, and it ends up detecting only the metal bulk and fluctuations of the surface. All other explored spatial resolutions (red circles in Fig. 4E) are shown in supplementary Figure S3.
@@
Figure 5: Optimal resolution and characteristic length-scales in experimental complex systems: e.g., collective waves in Quincke rollers colloids. A Cartoon representation of Quincke rollers: dielectric colloidal microparticles confined in the $xy$ plane that, immersed in a conducting fluid, give rise to complex collective motions, waves, vortexes, etc., under exposure to a weak DC electric field (orthogonal to the $xy$ plane). We analyze trajectories resolved from an experimental microscopy movie taken from Ref. [liu_activity_2021], where 6921 Quincke rollers undergo a collective wave that crosses a microscopy field of $700\times 700$ $\mu$m2 left-to-right along $\tau=0.25$ s of observation. We measured the average local velocity alignment using a sphere of radius $r_{c}$ (Eq. 1). B Radial distribution function $g(r)$ of the particles in the system, and characteristic $r_{c}$ distances identified by red circles. C Onion plot for the analysis of the $\phi$ time-series computed with cutoff radius $r_{c}=58.8$ $\mu$m: number of clusters (in blue) and fraction of unclassifiable data (orange) as a function of the time-resolution $\Delta t$ (the red dashed line identifies the $\Delta t=2$ ms time-resolution, for which results are shown as an example in panels D-E). D The five clusters resolved by Onion Clustering at the example time-resolution, colored in gray, blue, yellow, green, and red (ordered with increasing $\phi$). In red is clearly visible the core of the wave crossing the microscopy field (see also Supplementary Movie S1). E Four snapshots taken from the trajectory, where the Quincke rollers are colored according to the cluster they belong to. F-H Onion results of the analysis of the $\phi$ time-series calculated with different $r_{c}$ values. In this case, the maximum efficiency in information extraction-and-classification is encountered at $r_{c}=91$ $\mu$m, corresponding to the $\sim 5th$ neighbors shell. I Number of clusters resolved in each system, rescaled so that the maximum of each curve is equal to 1, as a function of the cutoff radius $r_{c}$ used for the computation of the descriptor. In order to compare the different systems, $r_{c}$ is expressed in multiples of the first neighbors shell radius $a$.
@@
As a final test case, we focus on a recently reported active matter system, composed by mesoscopic polystyrene particles, known as Quincke rollers. Under certain conditions, these particles exhibit highly non-trivial collective dynamics, such as emerging waves, vortexes, etc [liu_activity_2021] (Fig. 5A). For specific details about the system and how the analyzed trajectories were obtained, we refer the interested reader to Refs. [liu_activity_2021, becchi_layer-by-layer_2024]. In particular, the analyzed trajectory is reconstructed from a microscopy movie, with a microscopic field of $700\times 700$ $\mu$m2, containing 6921 polystyrene particles (with a diameter of $\sim 10$ $\mu$m), which show a wave span a large area of the system from left to right. The particles’ positions in the $xy$ plain when exposed to a weak perpendicular electric field were tracked for a duration of approximately 0.25 s.
@@
Also for this mesoscopic system we perform an analysis similar to those presented in previous cases. Fig. 5B, shows the $g(r)$ computed from the trajectory, identifying the critical distances, marked by red circles (such distances characterize typical neighbors shells within the system, as seen previously). In this case, we employed a physic-inspired descriptor widely used for studying active matter  [ginelli_physics_2016, aung_local_2024], local velocity alignment $\phi$. Also for this system, we denoised the descriptor using local averaging. The resulting quantity measures the average local velocity alignment in the environment surrounding each particle. This provides another demonstrative case of how our analysis is transferable to the study of time-series obtained using any type of descriptor.
@@
The Onion Clustering analysis of the de-noised local velocity alignment time-series detects several statistically distinct environments in this system. Fig. 5C-D shows the results produced by Onion Clustering using a cutoff of 59 $\mu$m (Including up to the third neighbor shell, Fig. 5B) with a time resolution of $1.6$ ms (vertical red dashed line, Fig. 5C). With this spatial resolution, five clusters can be distinguished. The gray cluster includes the majority of the particles, having $\Phi_{i}\sim 0$, which are not moving or with small random relative motions. The red cluster includes the particles with the highest values of $\Phi_{i}$, located in the inner part of the wave, moving with the maximum alignment in a coordinate fashion. The three remaining clusters (green, yellow, blue) correspond to intermediate alignments velocities area, from highest to lowest values respectively (Fig. 5D). Looking at Fig. 5E, it is possible to refer such clusters to well-defined physical regions in the system ,where the red area corresponds to the center of the wave crossing left-to-right the microscopic field over the observation, while the blue,yellow and green areas represent the wave edges (see also supporting Movie1).
@@
The value $r_{c}=91$ $\mu$m, corresponding to the fifth neighbors shell, is found to be the optimal spatial resolution for resolving the internal complexity of this system. In fact, such characteristic length scale correspond to the width of the collective wave traversing the system (Fig. 5E:in red). Notably, using $r_{c}\geq 91$ $\mu$m means using a cutoff larger that the wave thickness which results in reducing the efficiency of the analysis: using $r_{c}\sim 112$ $\mu$m is found detrimental, and the analysis resolve again at most 4 distinct clusters (Fig. 5H). Essentially, this analysis reveals the characteristic size of the dynamic physical events captured by the descriptor within the system. All other explored spatial resolutions (red circles in Fig. 5B) are shown in supplementary Figure S4.

### v4

@@ Results
It is well-known that complex systems can be governed by local or collective events Cioni et al. (2023); Crippa et al. (2023a); Caruso et al. (2023); de Marco et al. (2021); Butail, Mwaffo, and Porfiri (2016); Offei-Danso, Hassanali, and Rodriguez (2022) that may emerge and amplify within them. When one does not have prior knowledge of a system, choosing the best analysis setup to capture the effects/phenomena characterizing its behavior may be non-trivial. Nonetheless, such a choice can influence or bias the analysis, potentially revealing or obscuring distinct types of information, depending on the underlying nature and physics of the system (Fig. 1F), as we are going to discuss in the subsequent analyses.
In the case of aqueous systems, it is known that focusing solely on the first solvation shell can be limiting Crippa et al. (2023a); Offei-Danso, Hassanali, and Rodriguez (2022). It has been recently demonstrated how analyses extending beyond that (e.g., including up to the fifth nearest neighbor)Martino et al. (2024), can offer a more comprehensive understanding of the system Donkor, Laio, and Hassanali (2023); Donkor et al. (2024); Foffi and Sciortino (2023); Soper (2019); Daidone et al. (2023). This raises important questions such as, e.g.: What is the effect of the spatial resolution on the results that can be effectively attained upon analysis? Is there an “optimal resolution" to analyze a system? And is it possible to somehow determine it a priori? Addressing such questions is challenging, but at the same time particularly useful, especially in the study of systems about which little is known a priori.
Fig. 1 shows the results of analyses obtained calculating the LENS time series using a cutoff of $r_{c}=10$. While an $r_{c}$ around such value is commonly assumed to enable information-rich analyses of aqueous systems, Crippa et al. (2023a); Martino et al. (2024); Caruso et al. (2023); Donkor, Laio, and Hassanali (2023); Capelli, Muniz-Miranda, and Pavan (2022) we introduce an automated approach to systematically determine the optimal spatial resolution to study virtually any type of system. It is important to emphasize that the extractable information – and the corresponding optimal cutoff – depends both the internal physics of the system and the descriptor’s effectiveness in capturing relevant physical features. The efficiency and signal-to-noise ratio are in fact intrinsic properties of each descriptor Martino et al. (2024). Here, using LENS as an example, we present a systematic, data-driven approach to determine the optimal resolution for studying the internal dynamic complexity and overall physics of the system.
@@
We calculated the LENS descriptor for all oxygen atoms throughout the trajectory, using the identified key distances as cutoffs. For each cutoff, our method applied Onion Clustering to analyze the time series and determine the number of statistically distinct environments detected. By gathering and comparing these results, we identified the “optimal spatiotemporal resolution" for extracting physically relevant information from the trajectories. At this stage, the analysis depends solely on the chosen descriptor.
@@
Fig. 2B-G show the results obtained via Onion Clustering of the LENS time series computed with the different cutoffs. Fig. 2B illustrates the results of Onion Clustering applied to LENS time series computed with $r_{c}=3.3$, corresponding to the first solvation shell. At all time resolutions, the analysis detected at most one dynamically persistent environment, the bulk (ice), while all other molecular data points were grouped into the unclassifiable data cluster (transitions). Even at the highest resolution, the fraction of unclassifiable data points was approximately 50%, highlighting the limited effectiveness of considering only the first solvation shell for this aqueous system. This outcome is akin to conventional pattern recognition methods. Moreover, for time resolutions below $\Delta t>10$ ns, Onion Clustering fails to resolve any dynamically persistent clusters (Fig. 2B), as the fraction of unclassifiable data points (orange) reaches 100%. These results show that restricting the analysis to the first solvation shell does not includes in the LENS descriptor enough information for a meaningful classification of the system’s dynamics.
@@
Notably, when the second solvation shell is included using a cutoff of $r_{c}=7.5$ (Fig. 2C), Onion Clustering identifies 2 to 3 dynamically persistent environments. Crucially, the ice/water interface becomes discernible within this time series for resolutions between 2 and 8 ns. This observation aligns with findings reported for other descriptors Martino et al. (2024); Lionello et al. (2024), which highlight the existence of a lower time resolution threshold necessary to identify certain environments. In fact, for $\Delta t>8$ ns, the analysis loses the ability to distinguish the ice/water interface as a separate environment. Interestingly, Fig. 2C also shows that at very high time resolutions ($\Delta t<2$ ns), information about the ice/water interface is again lost, reducing the number of detectable clusters to two. This, as recently discussed  Martino et al. (2024); Lionello et al. (2024), is due to the fact that there is a minimum observation time which is required so that transitions can be robustly distinguished from the more dominant states, and thus classified as separate environments.
@@
When the cutoff distance is increased beyond 8.5 Å (Fig. 2D-G), including the third solvation shell, the LENS time series consistently resolve three dynamically persistent clusters down to a resolution of $\Delta t\sim 10$ ns. This clustering stability persisted up to a cutoff distance of approximately 15 Å. In other words, spatial resolutions in the range of $8.5$to 15 Åoffer the best performance, as they maximize the detection of distinct clusters (Fig. 2D-F) and minimize information loss (orange curve) for time resolutions $\Delta t<10$ ns. For this system, these spatial resolutions, when applied to the analysis of LENS time series, enable the extraction and representation of the system’s dominant features with optimal efficiency.
@@
Figure 3: Effect of spatial and temporal resolutions on information extraction from, e.g., SOAP time series data. A Schematic representation of SOAP. For each molecule $i$ in the system, at each sampled timestep $t$, the SOAP vector contains information on the distances and spatial displacements of the neighbors within a sphere of cutoff $r_{c}$: the SOAP power spectrum is thus a fingerprint of the local neighbors density around every SOAP center (the oxygen atoms of each water molecule, in this case). B-G Onion plots reporting the results of the clustering analyses of the SOAP PC1 time series as a function of the cutoff $r_{c}$ values used to calculate the SOAP power spectra. H In green: mean number of classifiable micro-clusters (ENVs) before the fraction of unclassifiable data reaches 50%, as a function of the $r_{c}$ used in the analysis. The data show a clear trend (green curve), where the maximum efficiency in information extraction-and-discretization is obtained at $13\leq r_{c}\leq 15$, where the interface is detected in a robust way down to time-resolutions of $\Delta t<20$ ns.
@@
When the LENS cutoff radius was set to values beyond $15$, the effectiveness of the analysis began to diminish. We extended the tests up to the maximum cutoff of 30Å, constrained by the size of the simulation box. As shown in Fig. 2B-G and in supplementary Fig. S1, the ability to distinguish the ice/water interface progressively deteriorates with increasing cutoffs. At higher values, the analysis yielded results similar to those obtained with smaller cutoffs, where only the solid ice phase can be classified as a stable environment.
@@
This result provides critical insights into the physics of the system, revealing the characteristic scales of the environments that dominate its dynamics. In this case, this ice/water system exhibits internal dynamics driven by collective events with specific spatial and temporal scales. For instance, the ice/water interface has a characteristic thickness of around 1.5 nm and contains molecules with a residence time of up to $\sim 10$ ns (Fig. 2I). Clearly, this “optimal spatiotemporal resolution" depends on the dynamics of each system (see next paragraph) and on the capability of different descriptors to capture them. Each descriptor emphasizes distinct features and has its own signal-to-noise ratio Martino et al. (2024). To illustrate the generality of these concepts, we repeated the Onion Clustering analyses on the first principal component (PC1) time series of the SOAP spectra computed for each molecule, calculated with different cutoff radii as done for LENS. This serves as another example of a descriptor providing rich information on the system (see Method for details).
@@
The results shown in Fig. 3 show a very similar trend as those in Fig. 2. As in the LENS analysis, Fig. 3A-G demonstrates the existence of an optimal spatial resolution that maximizes the information extracted from the SOAP PC1 time series. Below this resolution, the analysis lacks sensitivity, and above it, the time series become too noisy to resolve the ice/water interface. All other explored spatial resolutions are shown in supplementary Fig. S2. Also MD representative snapshot of this system, all taken at $\Delta t=40$ ns are shown in supplementary Fig. S5. The plot in Fig. 3H shows a trend similar to Fig. 2H, revealing an optimal spatial resolution at $r_{c}=13-15$. With this $r_{c}$, the interface is detectable using a time-resolution $\Delta t<20$ ns (Fig. 3D-E). Interestingly, the slightly higher optimal cutoff radius for SOAP compared to LENS reflects differences between the descriptors: while LENS captures exchanges between molecules inside and outside the cutoff, SOAP considers only the local environment within the cutoff. This makes the SOAP optimal radius consistent with the detection of events at the fourth solvation shell, compared to the third shell for LENS, while capturing the same underlying physical phenomena. Let’s now apply these concepts to another system with completely different dynamical complexity to investigate how the approach performs in distinct scenarios.
@@
Figure 4: Optimal resolution in the study of a dynamic metal surface. A MD snapshots of a Cu surface composed of 2400 atoms simulated at $T=600$ K: the atoms are colored according to their coordination number. Top: crystalline surface at $T=0$ K before simulation start; bottom: representative snapshot of the surface equilibrated at $T=600$ K. B Onion plot: number of clusters (in blue) classified from the LENS time series obtained with cutoff radius $r_{c}=5$, and fraction of unclassifiable data points (orange) as a function of the time-resolution $\Delta t$ used in the analysis. The vertical red dashed line indicates $\Delta t=0.12$ ns as an example time-resolution used to plot the analysis results in panels C-D. C LENS time series for all Cu atoms in the surface model along the entire MD simulation of $\tau=150$ ns (sampled every $\Delta\tau=10$ ps - see Methods for complete details). At the resolution of $\Delta t=0.12$ ns, Onion Clustering classifies six different micro-clusters (colored areas) whose thresholds are identified by horizontal colored dashed lines. D Top view of the simulation box at three different simulation times: $t=40$, $61$, and $150$ ns. An example Cu atom (ID = 144) is highlighted in black, whose history trajectory up to that point is colored according to the LENS micro-cluster it belonged to. At 150 ns, a red/orange straight vertical sliding on the (211) surface edge is clearly visible. E Radial distribution function $g(r)$ of the Cu atoms in the system: relevant spatial cutoffs are highlighted by red circles (relevant $g(r)$ minima). F-M Onion plots reporting the results of the analyses on LENS time series with different $r_{c}$. Same color coding as before. N Mean number of clusters (ENVs) classifiable (before ENV0 reaches 50%) as a function of the $r_{c}$ used in the analysis. In this case, dominated by local single-atom dynamical events, the optimal resolution is achieved at the smallest $r_{c}$: i.e., when accounting only for the first neighbors shell.
@@
To demonstrate the generality of our approach, we applied it to analyze LENS time series computed from simulation of an atomistic model of a copper (211) surface simulated at 600 K (Fig. 4A) composed by 2400 Cu atoms replicating on the $xy$ plane via periodic boundary conditions, using a deep neural network potential trained to guarantee DFT-level accuracy (see Method section for details) Cioni et al. (2023). As recently demonstrated, even well below the melting temperature $\sim 1400$ K, such system demonstrates a peculiar dynamics, with sparse individual atoms migrating along the surface crests and diffusing sliding rapidly over the surfaces face, being then reincorporated into the crests Cioni et al. (2023); Crippa et al. (2023a, b); Caruso et al. (2024). These rare local events are challenging to capture using conventional pattern recognition method, while they can be effectively captured via Onion Clustering analyses of LENS time series Becchi, Fantolino, and Pavan (2024).
Using this MD trajectory, we followed the same procedure illustrated in the previous paragraph, computing LENS using different cutoff radii (using different $g(r)$ minima, Fig. 4E) and then measuring, via Onion Clustering, how much information can be extracted for each time series dataset. As shown in Fig. 4, Onion Clustering can detect different environments. For every choice of the cutoff radius, the clustering can identify various, more or less dynamic, environments in the system (Fig. 4B-C, for instance, using $r_{c}=5$) corresponding, e.g., to the metal bulk or static surface domains, as well as the crests and diffusing atoms on the surface (Fig. 4D). The trend in the number of resolvable clusters (Fig. 4F-M: in blue) shows how the number of distinct environments that Onion Clustering can resolve in the system decreases as the cutoff radius used in the analysis increases (Fig. 3B,F-M). In Fig. 4F, for example, it is evident that up to 8-10 clusters can be resolved using $r_{c}=3$(first neighbors shell) up to a temporal resolution of $\Delta t\sim 100$ ps (the faster dynamics occurring in this system require an higher temporal resolution than the ice/water system studied previously). As the cutoff increases, the number of clusters tends to decrease, reaching a maximum of 8 for a cutoff radius in the range $6.5\leq r_{c}\leq 10.5$. For $r_{c}>14$, there is a considerable drop in the number of resolved clusters, and above $r_{c}=15$the effectiveness of the analysis deteriorates, and it ends up detecting only the metal bulk and fluctuations of the surface. All other explored spatial resolutions (red circles in Fig. 4E) are shown in supplementary Fig. S3.
@@
Figure 5: Optimal resolution and characteristic length-scales in experimental complex systems: e.g., collective waves in Quincke rollers colloids. A Cartoon representation of Quincke rollers: dielectric colloidal microparticles confined in the $xy$ plane that, immersed in a conducting fluid, give rise to complex collective motions, waves, vortexes, etc., under exposure to a weak DC electric field (orthogonal to the $xy$ plane). We analyze trajectories resolved from an experimental microscopy movie taken from Ref. Liu et al. (2021), where 6921 Quincke rollers undergo a collective wave that crosses a microscopy field of $700\times 700$ $\mu$m2 left-to-right along $\tau=0.25$ s of observation. We measured the average local velocity alignment using a sphere of radius $r_{c}$ (Eq. 1). B Radial distribution function $g(r)$ of the particles in the system, and characteristic $r_{c}$ distances identified by red circles. C Onion plot for the analysis of the $\phi$ time series computed with cutoff radius $r_{c}=58.8$ $\mu$m: number of clusters (in blue) and fraction of unclassifiable data (orange) as a function of the time-resolution $\Delta t$ (the red dashed line identifies the $\Delta t=2$ ms time-resolution, for which results are shown as an example in panels D-E). D The five clusters resolved by Onion Clustering at the example time-resolution, colored in gray, blue, yellow, green, and red (ordered with increasing $\phi$). In red is clearly visible the core of the wave crossing the microscopy field (see also Supplementary Movie S1). E Four snapshots taken from the trajectory, where the Quincke rollers are colored according to the cluster they belong to. F-H Onion results of the analysis of the $\phi$ time series calculated with different $r_{c}$ values. In this case, the maximum efficiency in information extraction-and-classification is encountered at $r_{c}=91$ $\mu$m, corresponding to the $\sim 5th$ neighbors shell. I Number of clusters resolved in each system, rescaled so that the maximum of each curve is equal to 1, as a function of the cutoff radius $r_{c}$ used for the computation of the descriptor. In order to compare the different systems, $r_{c}$ is expressed in multiples of the first neighbors shell radius $a$.
@@
As a final test case, we focused on a recently reported active matter system, composed by mesoscopic polystyrene particles, known as Quincke rollers. Under certain conditions, these particles exhibit highly non-trivial collective dynamics, such as emerging waves, vortexes, etc Liu et al. (2021) (Fig. 5A). For specific details about the system and how the analyzed trajectories were obtained, we refer the interested reader to Refs. Liu et al. (2021); Becchi, Fantolino, and Pavan (2024). In particular, the analyzed trajectory is reconstructed from a microscopy movie, with a microscopic field of $700\times 700$ $\mu$m2, containing 6921 polystyrene particles (with a diameter of $\sim 10$ $\mu$m), which showed a wave span a large area of the system from left to right. The particles’ positions in the $xy$ plain when exposed to a weak perpendicular electric field were tracked for a duration of approximately 0.25 s.
@@
Also for this mesoscopic system we performed an analysis similar to those presented in previous cases. Fig. 5B, shows the $g(r)$ computed from the trajectory, identifying the critical distances, marked by red circles (such distances characterize typical neighbors shells within the system, as seen previously). In this case, we employed a physic-inspired descriptor widely used for studying active matter  Ginelli (2016); Aung, McClure, and Abaid (2024), local velocity alignment $\phi$. Also for this system, we denoised the descriptor using local averaging. The resulting quantity measures the average local velocity alignment in the environment surrounding each particle. This provides another demonstrative case of how our analysis is transferable to the study of time series obtained using any type of descriptor.
@@
The Onion Clustering analysis of the de-noised local velocity alignment time series detects several statistically distinct environments in this system. Fig. 5C-D shows the results produced by Onion Clustering using a cutoff of 59 $\mu$m (Including up to the third neighbor shell, Fig. 5B) with a time resolution of $1.6$ ms (vertical red dashed line, Fig. 5C). With this spatial resolution, five clusters can be distinguished. The gray cluster includes the majority of the particles, having $\Phi_{i}\sim 0$, which are not moving or with small random relative motions. The red cluster includes the particles with the highest values of $\Phi_{i}$, located in the inner part of the wave, moving with the maximum alignment in a coordinate fashion. The three remaining clusters (green, yellow, blue) correspond to intermediate alignments velocities area, from highest to lowest values respectively (Fig. 5D). Looking at Fig. 5E, it is possible to refer such clusters to well-defined physical regions in the system ,where the red area corresponds to the center of the wave crossing left-to-right the microscopic field over the observation, while the blue,yellow and green areas represent the wave edges (see also supporting Movie1).
@@
The value $r_{c}=91$ $\mu$m, corresponding to the fifth neighbors shell, is found to be the optimal spatial resolution for resolving the internal complexity of this system. In fact, such characteristic length scale correspond to the width of the collective wave traversing the system (Fig. 5E:in red). Notably, using $r_{c}\geq 91$ $\mu$m means using a cutoff larger that the wave thickness which results in reducing the efficiency of the analysis: using $r_{c}\sim 112$ $\mu$m is found detrimental, and the analysis resolve again at most 4 distinct clusters (Fig. 5H). Essentially, this analysis reveals the characteristic size of the dynamic physical events captured by the descriptor within the system. All other explored spatial resolutions (red circles in Fig. 5B) are shown in supplementary Fig. S4.